A biophysicist seeks to separate subcellular particles with an analytic ultracentrifuge. The biophysicist must determine the magnitude of centripetal acceleration provided by the centrifuge of various speeds and radii.


Calculate the magnitude of centripetal acceleration at 8.4 cm from the centre of the centrifuge when it is spinning at 6.0 x 10^4 RPM. Find the answer in terms of g.


1
Expert's answer
2021-10-31T18:11:55-0400

"r=0.084 \;m \\

f = 60000 / 60 \;s \\

ac = 4 \pi^2 rf^2 \\

ac = 4 \times (3.14)^2 \times 0.084 \; m \times (\frac{60000}{60 \;s})^2 \\

= 3312825.6 \; m/s^2 \\

= 3312825.6 \times 0.1019716 \\

= 337814 \;g"